Gdub08:

write the slope intercept form of an equation of a line given two points: (-5,-1) , (-2,-3)

Gdub08:

I'm really stuck on this problem.

Gdub08:

The formula is y-y1=m(x-x1). To get m, u do (y2-y1)/(x2-x1)

neverknow:

y= -2/3x - 13/3 or y+1=-2/3 x ( x + 5) I think it's the second one though

Gdub08:

@neverknow wrote:
y= -2/3x - 13/3 or y+1=-2/3 x ( x + 5) I think it's the second one though
first one.

neverknow:

Ohh okay lol. I was confused cause you said said in Slope intercept and then said the formula was Point slope form. So that's my bad

Gdub08:

@neverknow wrote:
Ohh okay lol. I was confused cause you said said in Slope intercept and then said the formula was Point slope form. So that's my bad
ur good

neverknow:

Kk